OpenStudy (anonymous):

What is the solution set for |–2x + 16| – 9 = –3

OpenStudy (anonymous):

|–2x + 16|=6 now either -2x+16=-6 or -2x+16=6 Now solve

OpenStudy (anonymous):

these are the possible answers x = 5 and x = –11 x = –5 and x = –11 x = 5 and x = 11 x = –5 and x = 11

OpenStudy (anonymous):

x = 5 and x = 11